### v1

# First Astrometric Limits on Binary Planets and Exomoons orbiting Pictoris b

Abstract
The search for exomoons, or moons in other star systems, has attracted significant interest in recent years, driven both by advancements in detection sensitivity and by the expanding population of known exoplanets. The Pictoris system is a particularly favorable target, as its proximity and directly imaged planets allow for precise astrometric monitoring. We present astrometric constraints on the presence of binary planets and exomoons in the Pictoris system using archival observations from the GRAVITY interferometer and SPHERE instruments. We calculate these limits by modeling the motion of the two orbiting planets and introducing an additional perturbation to the model that simulates the astrometric motion caused by an exomoon orbiting the planet Pictoris b. We find that for short orbital periods ( days), a lunar companion is only allowed if its mass remains below () at confidence. At intermediate periods near 300 days, we exclude moons more massive than () at confidence. At longer orbital periods, we place the tightest constraints, ruling out any potential exomoon above () at days and () at days (both at confidence). These results place the first astrometric constraints on moons and binary planets in the Beta Pictoris system and demonstrate the sensitivity of interferometric observations for exomoon studies.
1 Introduction
As thousands of exoplanets have been discovered in the past 30 years, interest in whether moons orbit these worlds has grown. Despite the prevalence of moons in the Solar System (over 400 known111https://ssd.jpl.nasa.gov/sats/discovery.html) and the detection of a few candidates outside the Solar System (e.g. Teachey & Kipping, 2018; Lazzoni et al., 2020; Kipping et al., 2022), no extrasolar moons, or exomoons, have been confirmed. Nevertheless, given the well-established moon formation pathways benchmarked in the Solar System, theoretical studies suggest that exomoons should exist. One dominant mechanism in the Solar System is co-formation with the planets in circumplanetary disks, where moons form alongside their host planets in a gas-rich disk, similar to the Galilean moons of Jupiter (Canup & Ward, 2006). Another formation pathway is capture events, where a passing object becomes gravitationally bound to a planet. This process has been proposed as a key mechanism for the formation of moons on eccentric, inclined, and often retrograde orbits in our Solar System (Hamers & Portegies Zwart, 2018), including Triton (Agnor & Hamilton, 2006). Furthermore, debris generated from massive impacts, such as the hypothesized origin of the Earth-Moon system, could also lead to exomoon formation (Barr & Bruck Syal, 2017).
The hunt for exomoons is especially important for understanding the formation and evolution of planetary systems. Just as the Moon plays a key role in Earth’s orbital stability and tides, exomoons are expected to gravitationally influence their host planets, potentially affecting their rotational dynamics and long-term climate stability (Laskar et al., 1993; Poon et al., 2024). The successful detection of an exomoon also has the potential to transform our understanding of migration theories and lunar habitability (Barnes & O’Brien, 2002; Sucerquia et al., 2019, 2020; Martínez-Rodríguez et al., 2019).
Despite the importance of discovering exomoons, their detection remains extremely challenging. The signals from moons are often too subtle and faint to detect directly as a result of the limited resolution and sensitivity of current telescopes. Indirect methods, such as analyzing light curve variations during planetary transits, offer alternative ways to detect exomoons (Kipping et al., 2015; Teachey & Kipping, 2018; Kipping, 2021). Researchers search for signals such as transit timing variations (TTVs, Kipping, 2009, 2020a, 2020b), transit duration variations (TDVs, Kipping et al., 2009), and transit radius variations (TRVs, Rodenbeck et al., 2020). From these methods, a handful of exomoon candidates have been proposed, including Kepler-1625b-i, a potential Neptune-sized moon orbiting a super-Jovian planet (Teachey & Kipping, 2018), and Kepler-1708b-i, a potential 2.6 Earth-radii moon orbiting a Jupiter-sized planet at a distance of 12 planetary radii (Kipping et al., 2022). However, none of these candidates have been confirmed, and the detections are controversial (e.g. Kreidberg et al., 2019; Heller & Hippke, 2024). We note that similar methods can also be used to detect binary planets (Chakraborty & Kipping, 2022), but none have been identified.
With no confirmed exomoons, research has instead focused on placing increasingly strong limits on their possible presence. For instance, a considerable amount of work has gone into placing upper bounds on the masses and radii of exomoons based on transit observations (e.g. Kipping et al., 2013, 2014, 2015; Kipping, 2020a, b). More recently, other detection methods have been used to place constraints, including radial velocity (RV) observations of directly imaged planets (Vanderburg et al., 2018; Vanderburg & Rodriguez, 2021; Ruffio et al., 2023; Horstman et al., 2024). These findings demonstrate that, although the confident detection of exomoons is still elusive, current methods can place constraints on their masses and orbits.
Several studies have suggested that astrometry can be used to identify and constrain lunar companions within a system (Cabrera & Schneider, 2007; Agol et al., 2015; van Woerkom & Kleisioti, 2024). This is a particularly exciting method given the recent advancements in measuring precise astrometry of exoplanets using long-baseline interferometry. Recently, Lacour et al. (2021) showcased the high-precision astrometric capabilities of the GRAVITY instrument (GRAVITY Collaboration et al., 2017) on the Very Large Telescope Interferometer (VLTI), which can make astrometric measurements with precision of tens of microarcseconds. Building on these capabilities, Winterhalder et al. (2025) demonstrated through simulations and VLTI/GRAVITY observations that moons as small as 0.14 could already be inferred around the nearby (d26.9 pc) AF Lep b system, with future facilities expected to reach Earth-like moons in habitable zones. Complementary work by Wagner et al. (2025) presented a simulation framework for detecting exomoons around imaged giant planets, showing that Earth-mass moons in the habitable zone of Centauri A could be detected with future space- or ground-based facilities. Most recently, Kral et al. (2025) placed constraints on moons orbiting the brown dwarf HD 206893 B with GRAVITY astrometry and detected a tentative candidate. Together, these studies highlight how astrometry is emerging as a promising pathway toward the first confirmed detections of exomoons.
One of the most promising targets for astrometric studies is the Beta () Pictoris system, a young and nearby planetary system that has been extensively studied since the discovery of its circumstellar debris disk in the 1980s (Smith & Terrile, 1984). This system hosts two directly imaged giant planets, Pic b and c, making it ideal for testing astrometric detection techniques. Pic b, located at a semi-major axis of AU, was discovered through direct imaging, while Pic c, at AU, was discovered through RV measurements (Lagrange et al., 2010, 2020). Lagrange et al. (2020) combined astrometry and RV data to constrain the orbits and masses of both planets, finding that Pic b has a mass of and Pic c has a well-constrained mass of .
In this work, we use archival data from Lagrange et al. (2020) and Lacour et al. (2020a) to place constraints on the presence of exomoons and binary planets in the Pictoris system using high-precision astrometric measurements from the GRAVITY instrument. While we find no evidence supporting the existence of an exomoon, our analysis establishes strong constraints on the mass and orbital period of any potential companions, providing a framework for applying astrometric techniques to future exomoon searches. Our paper is organized as follows: Section 2 outlines the instrumentation and data. Section 3 describes our analysis methods, including the development and application of our model. Section 4 presents our results, focusing on the derived mass and orbital constraints. Section 5 then discusses the implications of our findings, comparing astrometry with complementary techniques and addressing the potential for future studies. We conclude in Section 6.
2 Observations
In this work, we used existing archival direct imaging and interferometric observations of Pictoris b and c. Pictoris has been observed for over a decade by numerous instruments, and the precision of astrometric measurements has increased over time. We focused on data taken with the GRAVITY interferometer (GRAVITY Collaboration et al., 2017) and the SPHERE high-contrast imager (Beuzit et al., 2019), provided in Table 1. Following Lacour et al. (2020a), we excluded earlier, less-precise data from the NACO (NAOS-CONICA) instrument (Lenzen et al., 2003; Rousset et al., 2003). Together, these datasets provide a high-precision foundation for our orbital analysis.
| Planet Name | Epoch (MJD) | RA (mas) | RA Err (mas) | DEC (mas) | DEC Err (mas) | Instrument | Correlation |
| Beta Pic b | 56999.23 | -188.44 | 4.4 | -294.74 | 3.99 | SPHERE | 0 |
| Beta Pic b | 57058.09 | -178.94 | 2.85 | -279.85 | 2.81 | SPHERE | 0 |
| Beta Pic b | 57058.09 | -180.52 | 1.17 | -282.71 | 1.17 | SPHERE | 0 |
| Beta Pic b | 57296.34 | -143.09 | 3.14 | -219.89 | 2.98 | SPHERE | 0 |
| Beta Pic b | 57296.34 | -144.9 | 1 | -217.92 | 1 | SPHERE | 0 |
| Beta Pic b | 57356.25 | -132.93 | 4.23 | -202.45 | 3.98 | SPHERE | 0 |
| Beta Pic b | 57356.25 | -134.32 | 1.52 | -202.08 | 1.52 | SPHERE | 0 |
| Beta Pic b | 57382.18 | -128.07 | 1.75 | -196.38 | 1.75 | SPHERE | 0 |
| Beta Pic b | 57382.18 | -130.96 | 3.35 | -195.71 | 3.31 | SPHERE | 0 |
| Beta Pic b | 57407.12 | -124.33 | 2.66 | -190.28 | 2.56 | SPHERE | 0 |
| Beta Pic b | 57407.12 | -125.1 | 1.73 | -187.57 | 1.73 | SPHERE | 0 |
| Beta Pic b | 57474.02 | -112.57 | 1.48 | -168.41 | 1.48 | SPHERE | 0 |
| Beta Pic b | 57474.02 | -113.69 | 2.62 | -168.9 | 2.55 | SPHERE | 0 |
| Beta Pic b | 57494 | -110.39 | 3.79 | -164.8 | 3.57 | SPHERE | 0 |
| Beta Pic b | 57494 | -106.79 | 1.7 | -161.71 | 1.7 | SPHERE | 0 |
| Beta Pic b | 57647.38 | -83.45 | 2.11 | -116.3 | 2.11 | SPHERE | 0 |
| Beta Pic b | 57647.38 | -79.507 | 3.81 | -116.28 | 3.76 | SPHERE | 0 |
| Beta Pic b | 57675.35 | -77.605 | 3.56 | -108.2 | 3.51 | SPHERE | 0 |
| Beta Pic b | 57675.36 | -76.58 | 2.54 | -111.22 | 2.54 | SPHERE | 0 |
| Beta Pic b | 57710.25 | -63.92 | 4.73 | -83.45 | 4.73 | SPHERE | 0 |
| Beta Pic b | 57710.27 | -77.255 | 10.7 | -105.18 | 10.29 | SPHERE | 0 |
| Beta Pic b | 58378.2 | 69.87 | 5.08 | 123.143 | 5 | SPHERE | 0 |
| Beta Pic b | 58378.35 | 68.54 | 2.92 | 127.13 | 2.92 | SPHERE | 0 |
| Beta Pic b | 58383.378 | 68.47 | 0.05 | 126.38 | 0.07 | GRAVITY | -0.86 |
| Beta Pic b | 58409.31 | 74.61 | 1.64 | 135.05 | 1.64 | SPHERE | 0 |
| Beta Pic b | 58409.38 | 77 | 3.57 | 138.44 | 4.04 | SPHERE | 0 |
| Beta Pic b | 58467.19 | 85.57 | 1.22 | 152.36 | 1.22 | SPHERE | 0 |
| Beta Pic b | 58467.2 | 85.93 | 4.12 | 154.926 | 3.75 | SPHERE | 0 |
| Beta Pic b | 58552.99 | 105.93 | 1.59 | 183.03 | 1.59 | SPHERE | 0 |
| Beta Pic b | 58790.3 | 142.55 | 2.16 | 248.71 | 2.27 | SPHERE | 0 |
| Beta Pic b | 58790.3 | 143.83 | 1.07 | 246.24 | 1.07 | SPHERE | 0 |
| Beta Pic b | 58796.17 | 145.514 | 0.108 | 248.589 | 0.045 | GRAVITY | -0.85 |
| Beta Pic b | 58798.356 | 145.65 | 0.03 | 249.21 | 0.09 | GRAVITY | -0.44 |
| Beta Pic b | 58839.19 | 153.56 | 2.42 | 258.71 | 2.57 | SPHERE | 0 |
| Beta Pic b | 58839.2 | 151.25 | 1.17 | 259.87 | 1.17 | SPHERE | 0 |
| Beta Pic b | 58855.065 | 155.41 | 0.15 | 264.33 | 0.29 | GRAVITY | -0.71 |
| Beta Pic b | 58887.07 | 159.91 | 0.94 | 270.94 | 0.94 | SPHERE | 0 |
| Beta Pic b | 58887.07 | 161 | 1.83 | 272.59 | 1.92 | SPHERE | 0 |
| Beta Pic b | 58889.139 | 160.96 | 0.06 | 273.41 | 0.13 | GRAVITY | -0.56 |
| Beta Pic b | 59221.238 | 211.59 | 0.02 | 352.62 | 0.05 | GRAVITY | -0.1 |
| Beta Pic b | 59453.395 | 240.63 | 0.09 | 397.89 | 0.04 | GRAVITY | -0.91 |
| Beta Pic c | 58889.14 | -67.36 | 0.17 | -112.59 | 0.24 | GRAVITY | -0.8 |
| Beta Pic c | 58891.065 | -67.67 | 0.11 | -113.2 | 0.19 | GRAVITY | -0.54 |
| Beta Pic c | 58916.043 | -71.88 | 0.07 | -119.6 | 0.14 | GRAVITY | -0.52 |
| Beta Pic c | 59220.163 | -52 | 0.2 | -80.86 | 0.34 | GRAVITY | -0.26 |
2.1 SPHERE
Observations of Pictoris b were acquired with the Spectro-Polarimetric High-contrast Exoplanet Research (SPHERE) instrument (Beuzit et al., 2019). SPHERE is equipped with the Infra-Red Dual-band Imager and Spectrograph (IRDIS, Dohlen et al., 2008) and the Integral Field Spectrograph (IFS, Claudi et al., 2008), enabling simultaneous high-contrast imaging. We make use of 34 SPHERE astrometric measurements from three different programs.
23 SPHERE measurements were obtained between December 2014 and September 2018 as part of the SHINE Guaranteed Time Observations survey and are reported in Lagrange et al. (2019). These observations provided the pre-conjunction orbital coverage of the planet and used the IRDIFS mode, with IRDIS operating in the H2 (central wavelength = 1.59 m; width = 0.055 m) and H3 (central wavelength = 1.67 m; width = 0.056 m) narrow bands, except in December 2014 when the K1 (central wavelength = 2.1025 m; width = 0.102 m) and K2 (central wavelength = 2.255 m; width = 0.109 m) filters were used. An additional 5 measurements were obtained after conjunction as part of the SHINE follow-up program and are reported in Lagrange et al. (2020). These observations were taken in October 2018 and December 2018 using the IRDIFS configuration (IRDIS H2/H3 + IFS in YJ mode, 0.95–1.35 m, ), and March 2019 using the IRDIS K1/K2 filters with the IFS in YH mode (0.95–1.65 m, ). For the March 2019 epoch, only IFS yielded a usable measurement due to IRDIS saturation. The final 6 observations of the system were obtained under a PI program led by Lagrange et al. (2020), all using the H2/H3 + YJ setup as in the SHINE observations. These measurements were taken in November 2019, December 2019, and February 2020.
All SPHERE observations followed established reduction techniques outlined in Chauvin et al. (2017) and Delorme et al. (2017). Specifically, the speckle calibration tool (SpeCal) developed for the SPHERE instrument by Galicher et al. (2018) was used in the reduction process to retrieve the position and brightness information of the identified point sources.
2.2 GRAVITY
The most important observations for our analysis (with the greatest ability to constrain exomoons) were taken with the GRAVITY instrument (GRAVITY Collaboration et al., 2017) on the Very Large Telescope Interferometer (VLTI) at ESO’s Paranal Observatory in Chile. GRAVITY collects K-band spectro-interferometric data and can achieve precise astrometric measurements with uncertainties ranging from 20150 as for directly imaged planets.
We use GRAVITY K band observations of Pictoris b and c from Lagrange et al. (2020) and Lacour et al. (2020a). The observations collected by Lagrange et al. (2020) took place on six nights between September 22, 2018, and February 9, 2020. Three of the observations were obtained in the medium-resolution (R = 500) mode of the GRAVITY spectrometer, and two were obtained in the high-resolution mode (R = 4000). On January 7th, both modes were used. Following Lacour et al. (2021), we exclude two of the GRAVITY observations of Lagrange et al. (2020) (MJDs 58738.267 and 58855.198) from our analysis due to poor coherence time and limited exposures. We also include in our analysis additional observations collected by Lacour et al. (2020a) as part of the ExoGRAVITY program (Lacour et al., 2020b) on the night of January 6, 2021 for Pictoris c and the nights of January 7 and August 27, 2021, for Pictoris b. The astrometric calibration from both Lagrange et al. (2020) and Lacour et al. (2020a) underwent processing using version 1.5.0 of the ESO GRAVITY pipeline (Lapeyrere et al., 2014).
3 Analysis
| Parameter | Two Planet Model | Two Planet & Moon Model |
| (M⊙) | ||
| Parallax (mas) | ||
| Error Scale Factor | ||
| SPHERE Offset RA (mas) | ||
| SPHERE Offset DEC (mas) | ||
| Orbital Parameters for Beta Pic b | ||
| Semi-Major Axis (au) | ||
| Eccentricity | ||
| cos(Inclination) | ||
| Argument of Periastron (deg) | ||
| Longitude of Ascending Node (deg) | ||
| Epoch of Periastron Passage | ||
| Orbital Parameters for Beta Pic c | ||
| Semi-Major Axis (au) | ||
| Eccentricity | ||
| cos(Inclination) | ||
| Argument of Periastron (deg) | ||
| Longitude of Ascending Node (deg) | ||
| Epoch of Periastron Passage | ||
| Orbital Parameters for Beta Pic b Moon | ||
| (M) | – | |
| Period (days) | – | |
| cos(Inclination) | – | |
| Longitude of Ascending Node (deg) | – | |
| Epoch of Periastron Passage | – | |
3.1 Brightness Ratio
We first investigated whether we must account for light emitted by any moon in our astrometric analysis. When we measure the astrometric wobble for two unresolved luminous sources, the motion of the “photocenter,” the center of flux of the two sources, will be smaller than the motion of either source. If a moon contributes significant flux compared to its host planet and we did not account for that in our analysis, we would infer an incorrect mass measurement or upper limit for that moon.
To evaluate the importance of this effect, we used the predicted K-band brightness of gas giant planets from ATMO evolutionary models (Phillips et al., 2020). For the moons, we took the K-band brightness of 23 Myr old gas giant planets with masses ranging from 0.5 to 10 . We then compared these values to the predicted K-band brightness of a 10 planet at that age, analogous to Pic b. The resulting brightness ratios are shown in Figure 1 as a function of the moon mass. Even for a 2 M moon, we predict a brightness ratio . This corresponds to a moon-planet mass ratio of , larger than any found within our own solar system. We therefore predict that any moons around Pic b will contribute negligible brightness, and that the photocenter’s motion will track the position of the planet. We proceed under the assumption that any exomoon is effectively dark, minimizing its impact on the measured astrometric motion.
3.2 Two Planet Model and Likelihood Function
To fit the SPHERE and GRAVITY data, we constructed a model capable of accurately predicting the positions of Pic b and c over time. To create this initial two-planet model, we adapted code from the Orbitize! Python package (Blunt et al., 2020) to simultaneously calculate the right ascension (RA) and declination (DEC) offsets of both planets with respect to the star. The model also accounts for the astrometric perturbation on Pic b caused by its inner companion, Pic c, following the methodology of Blunt et al. (2020).
We constructed a log likelihood function to compare the archival data (see Section 2) with this model. Maximizing the log likelihood () determines the best-fit orbital parameters for Pic b and c, and exploring the parameter space constrains the plausible parameter space for any orbiting exomoon. We found that the uncertainties on the GRAVITY observations appeared to be slightly underestimated, so we incorporated a multiplicative “jitter” term to allow our fit to adjust the GRAVITY uncertainties such that:
where is the multiplicative jitter factor, and represents uncertainties in either RA, or DEC for either planet b ( Pic b) or planet c ( Pic c).
The log likelihood consists of three components: the contributions from observations of Pic b and c ( and ), and Gaussian priors on various parameters, including the parallax and the masses and eccentricities of the planets. The final form of is expressed as:
where and are defined as:
In both Equations 3 and 4, the summations are taken over all observational astrometric epochs . and in Equation 3 refer to the observed astrometric offsets in right ascension and declination for Pic b, while and in Equation 4 refer to the same for Pic c. , , , correspond to the predicted positions of Pic b and c. and are the correlation coefficients for the astrometric measurements, and , , , and are the uncertainties in right ascension and declination for planet b and c . We note that the parameters in these equations are scaled by the “jitter” term described in Equation 1. We are not aware of any derivation of the normalized likelihood function for data points with correlated uncertainties (allowing us to fit for the best scaling factor for the correlated GRAVITY uncertainties), so we perform this calculation ourselves and describe it in detail in Appendix A.
We imposed informative Gaussian priors based on pulsation-corrected RV data from Lagrange (2019) on the masses of Pic b () and c (), as well as the eccentricity of Pic c (). Additionally, we imposed the same Gaussian prior on the star’s trigonometric parallax as Lacour et al. (2021) ( mas). For all other parameters, we used uniform priors with boundaries chosen either based on physical constraints (e.g., masses must be non-negative) or so that the posterior distributions for well-constrained parameters remain far from the boundaries. In total, our “Two Planet” model comprises a total of 19 free parameters. The full list of fitted parameters and priors we used is summarized in Table 2, and the best fit model is compared to observations in Figure 2.
3.3 Two Planet & Moon Model
To model the astrometric signature of a hypothetical dark exomoon orbiting Pic b, we extend the two-planet model described in Section 3.2 by adding a Keplerian companion on a bound orbit around Pic b. We used the same log likelihood functions as described in Equations 1 through 4, but include an additional astrometric perturbation term representing the motion of Pic b about the barycenter of the Pic b-moon system. To do this, we used orbitize to generate Keplerian position offsets (RA, DEC) for an exomoon orbiting Pic b, where its gravitational influence shifts the apparent position of Pic b by:
where and are the moon and planet masses, and and represent the astrometric offsets of the planet relative to the planet-moon center of mass (COM). These offsets are proportional to the moon’s position offsets () and scale by the ratio of the moon’s mass to the planet’s mass ().
We restricted our analysis to moons on circular orbits around Pic b and confined the moon’s orbital period to non-negative dynamically stable values based on the analysis of Domingos et al. (2006). In particular, the upper limit was derived based on the stability of prograde satellites, which must remain within:
where is the maximum prograde stable orbital radius, is the Hill sphere radius, is the planet’s orbital eccentricity, and is the exomoon’s orbital eccentricity (assumed to be zero). For Pic b, the prograde stability limit corresponds to an orbital radius of 0.532 AU and an orbital period of 1451.64 days. Based on these calculations, we impose a uniform prior on the exomoon’s orbital period with a lower bound defined by the Roche limit and an upper limit of 1500 days (which we truncated to the stability limit in post-processing). Overall, the addition of a potential exomoon requires five extra free parameters (moon mass, orbital period, inclination, position angle, and a reference time for the orbit), bringing the total number of fitted parameters from 19 in the two-planet model to 24 in the two-planet & moon model (see Table 2).
| Parameter | Two Planet System | Two Planet & Moon System | Lacour et al. (2021) |
| (M) | 10.02 | 10.03 | 10 |
| (M) | 9.030 | 9.030 | 9.15 |
| (M⊙) | 1.701 | 1.702 | 1.73 |
| Parallax (mas) | 51.39 | 51.39 | 51.44 |
| Error Scale Factor | 1.31 | 1.35 | – |
| SPHERE Offset RA (mas) | -0.10 | -0.09 | – |
| SPHERE Offset DEC (mas) | 0.06 | 0.05 | – |
| Orbital Parameters for Beta Pic b | |||
| Semi-Major Axis (au) | 10.03 | 10.04 | 9.95 |
| Eccentricity | 0.1094 | 0.1082 | 0.106 |
| Inclination (deg) | 88.991 | 88.994 | 88.99 |
| Argument of Periastron (deg) | 212.5 | 213.1 | 203.2 |
| Longitude of Ascending Node (deg) | 31.773 | 31.770 | 31.80 |
| Epoch of Periastron Passage | 0.774 | 0.775 | 0.730 |
| Orbital Parameters for Beta Pic c | |||
| Semi-Major Axis (au) | 2.688 | 2.681 | 2.61 |
| Eccentricity | 0.235 | 0.235 | 0.32 |
| Inclination (deg) | 88.94 | 88.94 | 88.92 |
| Argument of Periastron (deg) | 61 | 60 | 60.8 |
| Longitude of Ascending Node (deg) | 31.18 | 31.19 | 31.07 |
| Epoch of Periastron Passage | 0.825 | 0.823 | 0.708 |
3.4 Fitting
We fitted the data with our astrometric models in three steps to determine both best-fit parameters and uncertainties. First, we performed an exploratory MCMC analysis on our “Two Planet” model using the Python implementation of the differential evolution MCMC sampler (edmcmc, Vanderburg, 2021). This analysis successfully constrained the orbital parameters for the two-planet system, but was slow to achieve convergence due to the high-dimensional parameter space and the challenging covariances between different parameters.
Second, we used the best-fit values and posterior distributions for the orbital parameters to determine the prior ranges for a fit with pocoMC (Karamanis et al., 2022a, b), a Python implementation of a Preconditioned Monte Carlo algorithm. PocoMC uses machine learning to learn approximations of the posterior distributions that can substantially increase the efficiency of exploring parameter space. The result is a dramatic increase in the convergence speed for complex models like ours, particularly when incorporating the additional complexity of the exomoon parameters. We ran PocoMC on our “Two Planet” model using active particles and required at least effective particles per update step. Convergence was determined by monitoring the global effective sample size (ESS) and ensuring that it reaches a total of independent samples, indicating that the parameter chains are well-mixed and representative of the posterior distribution. We also examined the corner plot for each parameter, confirming sufficient sampling in the two-planet model. The model aligns well with the observed data (as shown in Figure 2), and our two-planet posterior distributions closely match those reported by Lacour et al. (2021). A full list of the best-fit parameters is given in Table 3.
Finally, we initialized a new PocoMC fit for our “Two Planet & Moon” model to establish upper limits for the mass of a putative moon orbiting Pic b. For this fit, we adopted the same configuration as in the “Two Planet” model, using active particles and effective samples. The final run consisted of a total of independent samples. The best-fit parameters for the “Two Planet & Moon” model are also given in Table 3.
4 Results
4.1 Constraints on Exomoons Around Pic b
First, we inspected the results of our astrometric fits to assess whether there is any evidence for an exomoon orbiting Pic b. When we compare the residuals of our best-fit “Two Planet” model to the results of our “Two Planet & Moon” model, as shown in Figure 3, we find that there are no significant remaining signals in the residuals of our two-planet fit. We also find that the posterior probability distribution for the mass of a moon in the “Two Planet & Moon” model is consistent with 0, indicating no significant signal of a moon in our dataset. We therefore proceed to investigate the upper limits on the mass of any putative exomoons orbiting Pic b from our “Two Planet & Moon” fits.
We show the two-dimensional posterior distribution of moon mass and orbital period, marginalized over all possible planetary and exomoon orbits, in Figure 4. In this figure, the color represents the posterior probability density, and we show corresponding , , and confidence limits as solid lines. We find that for short orbital periods ( days), our analysis places an upper limit on exomoon masses of approximately () with 95% confidence. At intermediate periods near days, we rule out companions exceeding (). At longer orbital periods, we place the tightest constraints on the moon mass: we constrain the maximum allowed mass to () at days and () at days. At the level, our constraints rule out exomoons more massive than () across most of the orbital period range. At some orbital periods around 300-500 days, we have lower sensitivity to exomoons, likely due to the relatively sparse observing cadence of the GRAVITY observations. Our , , and upper limits on the masses of companions to Pic b are listed in Table 4.
4.2 Orbital Parameters for Pic b and c
Our fits also allow us to investigate the sensitivity of the Pic orbital parameters to different assumptions and noise models. The results of our fits are summarized in Table 3. In general, our orbital parameters for Pic b and c are consistent with the values reported by Lacour et al. (2021). We also find good agreement between the orbital parameters determined by our “Two Planet” model and our “Two Planet & Moon” model, indicating that the presence of an unknown moon around Pic b would likely not bias our understanding of the planet’s orbit.
| Orbital Period (days) | 68% (M) | 95% (M) | 99.7% (M) |
| 2.424749 | 1.180788 | 3.009138 | 4.857555 |
| 7.274247 | 0.800393 | 1.896973 | 3.122645 |
| 12.123746 | 0.480739 | 1.014212 | 1.710106 |
| 16.973244 | 0.342458 | 0.675760 | 1.130730 |
| 21.822742 | 0.280340 | 0.543695 | 0.886558 |
5 Discussion
5.1 Constraints on Exomoons and Binary Planets in Context
In this work, we derive the first astrometric limits on exomoons and binary planets around Pic b. The limits that we achieve are driven by the ultra-precise astrometric measurements from the GRAVITY interferometer, which achieves 1-2 orders of magnitude better precision than traditional adaptive optics imaging. Our limits are strongest at long orbital periods, as expected due to the amplitude of astrometric perturbations scaling linearly with the semimajor axis of the companion. At the longest-period stable moon orbits, we can rule out moons more massive than about 30 Earth masses, or about twice the mass of Neptune.
The moons that we could detect in the Pic system are not like the moons in our own solar system. The Solar System’s most massive moon is Ganymede, which has a mass 50 times smaller than that of Earth, and 1500 times smaller than the lowest mass moon we can rule out at 3 around Pic b. However, the moons and binary planets that we could detect in the Pic system have similar mass ratios to some of the Solar System moons. At long orbital periods, we could detect moons with masses of about 1%, which is smaller than both the Pluto/Charon and Earth/Moon mass ratios. It is worth noting that both the Earth/Moon and Pluto/Charon systems are believed to have formed via collision and capture of lunar material (Denton et al., 2025; Canup & Asphaug, 2001); any moons we could detect might likely have formed similarly in their systems.
5.2 Are we sensitive to moons or binary planets?
Given that we are only sensitive to Neptune-mass companions and larger, it is worth asking: are the objects that we could detect moons, or should they instead be considered binary planets? The classification of a satellite as a moon or a binary planet remains an open question, particularly when considering systems beyond our solar system. A simple way to distinguish moons and binary planets would be to choose a mass ratio between the two orbiting objects above which an orbiting body is a binary planet, and below which it is a moon. Such a strategy has been applied to the IAU definition of a planet (Lecavelier des Etangs & Lissauer, 2022), which states that the mass ratio must be large enough that stable orbits can exist around the L4 and L5 Lagrange points: , or more precisely . If we were to, by extension, apply this definition to distinguish moons from binary planets, our analysis is sensitive to some exomoons around Pic b at orbital periods longer than about 100 days.
However, there is no official IAU definition for a moon, so instead, other criteria have been utilized. According to one commonly cited definition from Stern & Levison (2002), a moon is a satellite with a mass and semi-major axis small enough to keep the center of mass of the moon-planet system within the radius of the primary planet. In contrast, a binary planet is defined as a system in which the companion is sufficiently massive for the center of mass to extend beyond the primary planet. By this definition, we are not sensitive to exomoons around Pic b. We show the Stern & Levison (2002) boundary between moons and binary planets in Figure 4 as a solid green line. This boundary falls slightly below our 1 contour at all orbital periods and scales in a similar manner with orbital period to our sensitivity limits. This similar scaling is expected, since the Stern & Levison (2002) definition of a moon and the amplitude of astrometric signals both depend on the center of mass motion of the planet.
However, both of these definitions have drawbacks. For instance, although Charon formally meets the requirement for a binary planet from both of these two definitions, it is still commonly referred to as a moon of Pluto. Moreover, under the Stern & Levison (2002) definition, the same object at different orbital separations may be either a moon or a binary planet. Given orbital evolution, this implies that an orbiting companion may be both a moon and a binary planet at different times in the system’s history. For instance, while the Earth/Moon system currently meets the Stern & Levison (2002) criterion for a moon, after the Moon’s semimajor axis has increased by about 40% due to tidal evolution, it would by this definition formally become a binary planet. It is also plausible that because planet radii contract over time, a young planet/moon system may evolve into a binary system as the planet radius becomes small enough that it does not encompass the system’s center of mass. Ultimately, a precise distinction between moons and binary planets will likely have to wait until astronomers have a better understanding of the broader population beyond the solar system.
5.3 Prospects for Improved Exomoon Searches in the Future
As discussed in the preceding subsections, exomoon detection is still in its early stages, and we are not yet sensitive to Solar System-like moons around Pic b. It is therefore worth considering whether a pathway exists to detecting smaller moons around directly imaged planets using astrometry or other techniques.
There is already good potential for improvement to astrometric moon detection with existing facilities by collecting additional data. For example, our dataset only includes 7 GRAVITY observations of Pic b. Typical datasets used to search for planets have considerably more observations; for instance, Sahlmann et al. (2013) reported a planet candidate with 25 astrometric observations, and often planet discoveries using the radial velocity method have hundreds or thousands of observations (e.g. Feng et al., 2017). If more observations with GRAVITY were collected with the same observing strategy, we could plausibly expect the sensitivity to exomoons to improve by a factor of a few.
Another opportunity for improving sensitivity with astrometry will come from the planned GRAVITY+ upgrade to the GRAVITY interferometer. This upgrade will significantly enhance astrometric precision, making it possible to routinely achieve –as level uncertainties (Bourdarot et al., 2022). Compared to our existing dataset, which has a median uncertainty of 70 as and a mean uncertainty of 100 as, this upgrade could give an order-of-magnitude improvement in sensitivity. Since we are currently sensitive to moons, an order of magnitude improvement would give sensitivity to moons as low as a few Earth masses at long periods.
It will also be possible to probe a larger parameter space by combining astrometric observations with other techniques, in particular radial velocities (e.g. Vanderburg et al., 2018; Lazzoni et al., 2022). Even though both radial velocities and astrometry probe the “wobble” of the planet due to the moon, the signal amplitude of an astrometric orbit scales proportionally to the moon’s semimajor axis , while the amplitude of a radial velocity orbit scales with . Therefore, radial velocities are much more sensitive to short-period moons, while astrometry is more sensitive to long-period moons. Recently, Kenworthy et al. in prep reported the first constraints on exomoons around Pic b using radial velocity observations, achieving 350 precision and demonstrating sensitivity to sub-Saturn mass moons at short orbital periods. Combining these short-period radial velocity constraints with our strong astrometric constraints on long-period moons is a promising path forward to searching for moons at all orbital separations around directly imaged planets.
Both of these techniques could considerably improve as new telescopes are commissioned. Vanderburg et al. (2018) pointed out the opportunity to make extremely precise radial velocity observations of directly imaged planets with 30-meter class telescopes, where the improved angular resolution will dramatically improve the signal-to-noise ratio of planetary spectra. This could improve typical radial velocity precision from hundreds of meters per second to few precision like that achieved on stars today. On longer timescales, improving the capabilities of optical long-baseline interferometry could also aid in exomoon detection. For example, adding a fifth Unit Telescope to the VLTI (with a longer baseline) could considerably improve the angular resolution and astrometric precision of a successor to GRAVITY.
6 Conclusions
In this work, we used astrometric measurements from the GRAVITY interferometer and SPHERE imaging to place constraints on the presence of exomoons around Pic b. Our conclusions can be summarized as follows:
For short orbital periods ( days), we exclude moons more massive than () at the level.
At intermediate periods near days, we rule out companions exceeding () at the level.
At long orbital periods, we constrain the maximum allowed mass to () at days and () at days (both at the level).
These are the first astrometric limits on the presence of exomoons and binary planets in the Pictoris system, and to our knowledge, along with Kral et al. (2025), one of the first astrometric constraints on companions around any directly imaged planet. We find that existing observations are primarily sensitive to systems where the center of mass is outside of the planet itself, which is a commonly cited criterion for distinguishing binary planets from moons, (Stern & Levison, 2002). However, these observations are sufficient to achieve sensitivity to moon/planet mass ratios smaller than that of the Earth and its moon, and increasing the number of GRAVITY observations could plausibly extend sensitivity to companions that are unambiguously moons.
The derived astrometric limits highlight the power of astrometry in constraining exomoon orbital parameters, particularly when combined with interferometric measurements. Looking forward, synergy with future GRAVITY+ data and complementary RV measurements could further tighten constraints and enable confident detections of exomoons.
Data Availability
The percentile data used to generate Figure 4 are included in this paper’s supplementary material. An excerpt of the table is shown in Table 4.
VLTI, VLT:Melipal
Appendix A Scaling Factor and Normalization
This appendix details the derivation of the form of the normalization of a likelihood function that includes correlations between observations in the RA and DEC directions. Including the normalization term in the likelihood function we use in our astrometric fits allows us to adjust the scale of the correlated GRAVITY uncertainties within the fit.
A.1 Likelihood function
We define the likelihood contribution associated with observation to be:
where is a normalization constant and is an generalization of the traditional statistic to include correlated uncertainties. In particular, we define as:
As in Section 3.2, we define to be the correlation coefficient between astrometric measurements in the RA direction and the DEC direction for observation . Again, and refer to our astrometric measurements in those directions, while and refer to our model astrometric measurements in those directions.
As usual, to calculate the likelihood of all observations together, we take the products of their individual likelihood contributions:
A.2 Derivation of the Normalization constant
Because likelihood contributions are probabilities, they must integrate to 1 over all possible values of and :
We ensure this condition holds by evaluating the integral expression and solving for the normalization constant . We start by defining several constants and new variables for simplicity. In particular, we define:
These allow us to rewrite our expression for more compactly as:
We then define three more constants:
and substitute them into our expression for the likelihood contribution:
Returning to the normalization condition, we write:
By evaluating the multivariate Gaussian integral, we obtain the condition:
Solving for :
Substituting our original definitions for , , , and , we obtain: